### v1

# Distilling Machine Learning’s Added Value: Pareto Fronts in Atmospheric Applications

Abstract
While the added value of machine learning for weather and climate applications is measurable, explaining it remains challenging, especially for large deep learning models. Inspired by climate model hierarchies, we propose that a full hierarchy of Pareto-optimal models, defined within an appropriately determined error-complexity plane, can guide model development and help understand the models’ added value. We demonstrate the use of Pareto fronts in atmospheric physics through three sample applications, with hierarchies ranging from semi-empirical models with minimal tunable parameters (simplest) to deep learning algorithms (most complex). First, in cloud cover parameterization, we find that neural networks identify nonlinear relationships between cloud cover and its thermodynamic environment, and assimilate previously neglected features such as vertical gradients in relative humidity that improve the representation of low cloud cover. This added value is condensed into a ten-parameter equation that rivals the performance of deep learning models. Second, we establish a machine learning model hierarchy for emulating shortwave radiative transfer, distilling the importance of bidirectional vertical connectivity for accurately representing absorption and scattering, especially for multiple cloud layers. Third, we emphasize the importance of convective organization information when modeling the relationship between tropical precipitation and its surrounding environment. We discuss the added value of temporal memory when high-resolution spatial information is unavailable, with implications for precipitation parameterization. Therefore, by comparing data-driven models directly with existing schemes using Pareto optimality, we promote process understanding by hierarchically unveiling system complexity, with the hope of improving the trustworthiness of machine learning models in atmospheric applications.
With machine learning permeating the geosciences, it becomes urgent to distinguish incremental progress from enduring knowledge discovery. We show that Pareto-optimal hierarchies transparently distill the added value of new algorithms using three atmospheric physics applications, providing a timely complement to post-hoc explainable artificial intelligence tools.
1 Introduction
Have recent advancements in machine learning (ML) led to discoveries in atmospheric science? The added value of ML for weather and climate applications is measurable, but it often remains challenging to understand. Taking advancements in data-driven, medium-range weather forecasting (Ben-Bouallegue et al., 2023) as an example, increasing reliance on complex architectures makes state-of-the-art models difficult to interpret.
Weyn et al. (2019, 2020) used convolutional neural networks (CNN) with approximately 200k and 700k learned parameters to produce global forecasts that outperformed climatology, persistence, and a low-resolution numerical weather prediction (NWP) model for lead times smaller than one week. Following the success of early approaches, Rasp et al. (2020) developed a benchmark dataset for data-driven weather forecasting, which facilitated the objective assessment of rapid developments (e.g., Clare et al., 2021; Scher and Messori, 2021). Since then, for data-driven, medium-range forecasting, Rasp and Thuerey (2021) trained a M-parameter deep residual CNN, Keisler (2022) trained a M-parameter graph neural network (GNN), and Pathak et al. (2022) trained a M-parameter emulator combining transformers with Fourier neural operators. Recently, deep learning models started rivaling state-of-the-art, high-resolution, deterministic NWP models: Lam et al. (2022) via combined GNNs totaling M parameters, Bi et al. (2022) via a M-parameter Earth-like transformer, and Lang et al. (2024) via a M-parameter graph and transformer model. It is hard to pinpoint what makes these models so successful, even with modern explainable artificial intelligence (XAI; Buhrmester et al., 2021; Das and Rad, 2020) tools, given that XAI requires certain assumptions to be satisfied (Mamalakis et al., 2022, 2023) and involves choosing which samples to investigate, which is challenging for large models and datasets.
The growing complexity of data-driven models for weather applications shares similarities with the development of general circulation models (GCM) that followed the first comprehensive assessment of global climate change due to carbon dioxide (Charney et al., 1979). Unlike data-driven weather prediction, where reducing forecast errors could warrant increased complexity, GCMs have traditionally been created to not simply project but also comprehend climate changes (Held, 2005; Balaji et al., 2022). This implies that any additional complexity in an Earth system model should be well-justified, motivating climate model hierarchies that aim to connect our fundamental understanding with model prediction (e.g., Mansfield et al., 2023; Robertson and Ghil, 2000; Bony et al., 2013; Jeevanjee et al., 2017; Maher et al., 2019; Balaji, 2021).
Inspired by climate model hierarchies, we here show that modern optimization tools help systematically generate data-driven model hierarchies to model and understand climate processes for which we have reliable data. These hierarchies can: (1) guide the development of data-driven models that optimally balance simplicity and accuracy; and (2) unveil the role of each complexity unit, furthering process understanding by distilling the added value of ML for the atmospheric application of interest.
In this study, we showcase the advantages of considering a hierarchy of models with varying error and complexity, as opposed to focusing on a single fitted model (Fisher et al., 2019). After formulating Pareto-optimal model hierarchies (Sec. 2) and categorizing the added value of ML into four categories (Sec. 3), we apply our approach to three atmospheric processes relevant for weather and climate predictions (Sec. 4) to distill the added value of recently developed deep learning frameworks before concluding (Sec. 5).
2 Pareto-Optimal Model Hierarchies
In this section, we define Pareto optimality before discussing the definition of error and complexity.
2.1 Pareto Optimality
In multi-objective optimization, Pareto optimality represents a solutions set that cannot be improved upon in one criterion without worsening another criterion (e.g., Censor, 1977; Miettinen, 1999). The first step is to define a set of real-valued model evaluation metrics that we wish to minimize (e.g., error, complexity). We call a model Pareto-optimal w.r.t. these metrics and w.r.t. a model family if there is no model in that family that strictly outperforms in one metric while maintaining at least the same performance in all other metrics (e.g., Lin et al., 2019). The Pareto front (PF) is the set of all Pareto-optimal models, which can be defined using logical statements:
Intuitively, when we select a model from the Pareto front, any attempt to switch to a different model would mean sacrificing the quality of at least one evaluation metric. Conversely, a model that can be replaced without compromising any evaluation metrics is described as Pareto-dominated. In practice, we often seek to balance evaluation metrics measuring error and complexity, which we discuss in the next subsections.
2.2 Error
We emphasize the importance of holistic evaluation, which employs several error metrics with different behaviors: Traditional regression or classification metrics, distributional distances, spectral metrics, probabilistic scoring rules, reliability diagrams (Haynes et al., 2023), causal evaluation (Nowack et al., 2020), etc. To facilitate the use of Pareto-optimal hierarchies, we recommend prioritizing proper scores, whose expectation is optimal if and only if the model represents the true target distribution (Bröcker, 2009). For simplicity’s sake, we will employ mean squared error (MSE) as our primary error metric for our study’s applications. We make this choice because MSE is a proper score for deterministic predictions that can be efficiently optimized, while recognizing MSE’s inherent limitations for non-normally distributed targets.
2.3 Complexity
To our knowledge, there are no universally accepted metrics for quantifying the complexity of data-driven models within the geosciences. In statistical learning, various complexity metrics, such as Rademacher complexity (e.g., Bartlett et al., 2005), rely on dataset characteristics, whereas others, like the VC dimension (Vapnik and Chervonenkis, 2015), solely depend on algorithmic attributes. Here, we predominantly focus on two metrics that can be readily calculated from model attributes: the number of trainable parameters () and the number of features (). We choose these metrics due to their simplicity and versatility across the broad spectrum of models considered in our study. As we will confirm empirically in Sec. 4, and can be used as (very) approximate proxies for generalizability, as models with fewer trainable parameters and features with more stable distributions tend to result in superior generalizability. We defer the exploration of additional complexity metrics, such as the number of floating point operations (FLOP), to future work.
Equipped with these definitions, we can now ask: Why, along the Pareto front in a well defined error-complexity plane, does increasing complexity result in better performance? In the following section, we hence leverage Pareto optimality to distill machine learning’s added value.
3 The Distillable Value of Machine Learning
Amid rapid progress in optimization, machine learning architectures, and data handling, it can be challenging to distinguish long-lasting progress in modeling from small improvements in model error. We postulate that progress is more likely to be replicable if we can explain how added model complexity improves performance in simple terms. Based on this postulate, we propose simple definitions to categorize a model’s added value in the geosciences.
In geoscientific applications, we often work with features that are functions of space, discretized in spatial locations , and of time, discretized in timesteps . We will consider a model predicting a target vector from a multi-variate spatiotemporal field such that . We define a model as having added value w.r.t. a set of evaluation metrics and a simple baseline model if the following is true: When applied to representative out-of-sample data, is Pareto-optimal while the baseline model is not. Taking the data’s spatiotemporal structure into account, we organize a model’s added value into four mutually non-exclusive categories: functional representation, feature assimilation, spatial connectivity, and temporal connectivity (see Fig. 1).
3.1 Functional Representation
A fundamental aspect a model can improve is the functional representation of the observed relationship between a set of baseline features —fixed features used to benchmark performance—and the target . We deem to improve the functional representation of a baseline w.r.t. if and only if:
This improvement may stem from algorithms leading to better fits (e.g., gradient boosting instead of decision trees), improved optimization (e.g., the Adam optimizer and its variants instead of traditional stochastic gradient descent; Kingma and Ba, 2014), improved parsimony (e.g., by decreasing the number of trainable parameters via hyperparameter tuning) or enforced constraints (e.g., positive concentrations and precipitation). Improvements in functional representation are readily visualized via partial dependence plots or their variants (marginal plots to avoid unlikely data instances, accumulated local effects to account for feature correlation, etc.; Molnar, 2020). Note that Eq. 2 also captures improvements in probabilistic modeling by generalizing to a stochastic mapping and including probabilistic scores (Gneiting and Raftery, 2007; Haynes et al., 2023) in . From an atmospheric science perspective, improving functional representation helps identify nonlinear regimes, identify model extremes, and describe how sensitive the prediction is to different features. However, it can be challenging to faithfully describe these sensitivities if key features are missing from the baseline set of features .
3.2 Feature Assimilation
This motivates discussing the ability of a model to extract relevant information from a new feature set, which we refer to as feature assimilation and define as follows:
where we emphasize the difference between the baseline () and new () sets of features. This improvement may stem from previously unconsidered features (e.g., variables whose quality has increased in recent datasets), the ability of models to handle features whose information could not be extracted in past attempts (e.g., the ability of deep learning to extract nonlinear relationships), or the discovery of a more compact feature set that facilitates learning or improves generalizability. The assimilation of new features may improve interpretability (e.g, if some features simplify the target’s functional representation or if there are less features to consider), physical consistency (e.g., if the new feature set improves generalizability across regimes or consistency with physical laws), and predictability. While the features’ spatial and temporal discretizations could technically be considered part of feature selection, we choose to discuss them separately in the following subsections, given the central role of space and time in geoscientific applications.
3.3 Spatial connectivity
To formalize a model ’s improved ability to leverage the features’ spatial information, we distinguish the spatial locations used to discretize the baseline ’s features from the spatial locations used to discretize ’s features:
This improvement may stem from the ability to:
1.
handle features at different spatial resolutions (e.g., via improved pre-processing or handling of data). In atmospheric science, this can help consider multi-scale interaction and accommodate data from various Earth system models.
2.
hierarchically process spatially adjacent data (e.g., via convolutional layers). In atmospheric science, this acknowledges the high correlation between spatial neighbors due to, e.g., small-scale mixing.
3.
capture long-range spatial dependencies (e.g., via self-attention mechanisms or a graph structure), such as teleconnections in the Earth system.
3.4 Temporal connectivity
To formalize a model ’s improved ability to leverage the features’ temporal information, we distinguish the timesteps used to discretize the baseline ’s features from the timesteps used to discretize ’s features:
This improvement may stem from the ability to:
1.
handle features at different temporal resolutions (e.g., via improved pre-processing or handling of data). This can help consider multiple timescales and accommodate data from various Earth system models.
2.
process consecutive timesteps (e.g., via recurrent layers). This acknowledges the high temporal autocorrelation of Earth system data, which is a property of the underlying dynamical system.
3.
capture long-term temporal dependencies (e.g., via gating or self-attention mechanisms) and cyclic patterns (e.g., via data adjustments and temporal Fourier transforms) such as the diurnal and seasonal cycles.
4 Atmospheric Physics Application Cases
Building upon the four types of added value outlined in the previous section, this section demonstrates that Pareto optimality guides model development and improves process understanding through three realistic, atmospheric modeling case studies. Each case includes machine learning prototypes with demonstrated performance from previous studies, along with Pareto-optimal models newly trained for this study.
4.1 Cloud Cover Parameterization for Climate Modeling
4.1.1 Motivation
The incorrect representation of cloud processes in current Earth system models, with grid spacing of approximately 50-100 km (Arias et al., 2021), significantly contributes to structural uncertainty in long-term climate projections (Bony et al., 2015; Sherwood et al., 2014). Cloud cover parameterization, which maps environmental conditions at the grid scale to the fraction of the grid cell occupied by clouds, directly affects radiative transfer and microphysical conversion rates, influencing the model’s energy balance and water species concentrations.
Storm-resolving simulations with grid spacing below km reduce uncertainty related to the interaction between storms and planetary-scale dynamics by explicitly simulating deep convection (Stevens et al., 2020). However, their large computational cost prohibits their routine use for ensemble projections (Schneider et al., 2017). Machine learning can learn the storm-scale behavior of clouds from short, storm-resolving simulations, potentially improving coarser Earth system models through data-driven parameterizations (Gentine et al., 2021).
In this section, we aim to understand the improvement gained from the higher-fidelity representation of storms and clouds. As illustrated in Fig. 2, we demonstrate that this knowledge can be symbolically distilled into an analytic equation that rivals the performance of deep learning.
4.1.2 Setup
We follow the setup described in Grundner et al. (2024), to which readers are referred for details. Fields are coarse-grained from storm-resolving ICON simulations (Giorgetta et al., 2018) conducted as part of the DYAMOND inter-comparison project (Stevens et al., 2019; Duras et al., 2021). The original simulations use a horizontal grid spacing of km and 58 vertical layers below 21 km (the maximum altitude with clouds in the dataset). They are coarse-grained to a typical climate model resolution of km horizontally and 27 vertical layers, converting the binarized, high-resolution condensate field (1 if the cloud condensate mixing ratio exceeds kg/kg, 0 otherwise) into a fractional area cloud cover (unitless).
To prevent strong correlations between the training and validation sets, the union of the “DYAMOND Summer” (Aug 10 to Sep 10, 2016) and “DYAMOND Winter” (Jan 30 to Feb 29, 2020) datasets was partitioned into six consecutive temporal segments. The second segment (approximately Aug 21 to Sep 1, 2016) and the fifth segment (approximately Feb 9–19, 2020) form the validation set. For all models, excluding traditional methods, the features are standardized to have a mean of zero and a standard deviation of one within the training set.
Once coarse-grained and pre-processed, we aim to map the environmental conditions on vertical levels indexed by the background terrain-following height grid , to the cloud cover on the same vertical levels. The variables include the horizontal wind speed [m/s], specific humidity [kg/kg], liquid water mixing ratio [kg/kg], ice mixing ratio [kg/kg], temperature [K], pressure [Pa], and relative humidity . Except for the “non-local NN” in Sec. 4.4.1.4.1.4, we simplify the mapping to a “vertically quasi-local” one where cloud cover at a given level depends only on the atmospheric variables at the same level and their first and second-order derivative with respect to . also includes geometric height [m] and surface variables: land fraction [%] and surface pressure [Pa]. In summary, we approximate some mapping:
using a hierarchy of machine learning models. In the following subsections, we will show that Pareto-optimal hierarchies not only facilitate data-driven model development, allowing for the comparison of simple baselines with neural networks, but also promote process understanding when tracing the Pareto front.
4.1.3 Existing Baselines and Polynomial Regression
We start with the Sundqvist baseline (red star in Fig. 2; Sundqvist et al., 1989), the standard cloud cover parameterization in ICON. The Sundqvist parameterization implemented in ICON represents cloud cover as a monotonically increasing function of relative humidity, provided it exceeds a critical threshold (Roeckner et al., 1996):
where our implementation includes 4 tunable parameters: , with values listed in appendix A. When , cloud cover is given by the model:
whose output is:
To account for marine stratocumulus (low) clouds (Mauritsen et al., 2019), we use two different sets of 4 parameters for land and sea using a land fraction threshold of 0.5. The Sundqvist scheme is parsimonious with only 8 trainable parameters, but it performs poorly against high-resolution data, with MSE values as large as despite having been re-tuned to our training set.
Hypothesizing that the Sundqvist scheme’s large error is due to its lack of features, we test the effect of adding features one-by-one. For that purpose, we apply forward sequential feature selection, which greedily adds features using a cross-validated score (here MSE), to a standard multiple linear regression model that includes polynomial combinations of all available features, up to a maximum degree of 3.
We find that linearly including temperature as a feature alongside relative humidity is enough to outperform the Sundqvist baseline, providing our simplest example of feature assimilation:
where is a two-feature linear regression whose output is:
with parameters , , and listed in appendix A. This feature assimilation result corresponds to a well known result: accurate cloud cover parameterization requires temperature in addition to relative humidity.
Adding more features to linear (red crosses), quadratic (orange crosses), and cubic (yellow crosses) polynomials further reduces error. However, incorporating a simple physical constraint:
i.e., that the absence of condensates implies no clouds, further improves the results. Note that this constraint can also be derived from data using a binary decision tree. Accounting for this constraint, polynomial models achieve MSEs less than half of the Sundqvist scheme’s (see Fig 2’s lower error, higher complexity polynomial models). Focusing on the Pareto-optimal, multilinear and quadratic model outputs:
with parameters listed in appendix A. Analyzing this first Pareto front unveils the role of each added feature in better representing cloud cover. First, it highlights the role of cloud condensates, both liquid and ice, in accurately describing cloud cover. Second, among all possible vertical gradients, Pareto-optimal models select the negative gradient of relative humidity to depict the role of inversions in increasing cloud cover, which is particularly important for low clouds, including marine stratocumulus clouds. However, upon closer inspection, these contributions are not always physically consistent: while the multilinear model correctly captures the increase of cloud cover with ice, it incorrectly decreases cloud cover when there is humidity above. Conversely, the quadratic model accurately captures the inversion role when temperature and humidity are fixed to their training set mean but wrongly assumes that inversions’ effect on cloudiness depends on humidity and temperature, and the model’s decrease in cloudiness with increasing mixed-phase cloud condensates suggests that its term is not an interpretable sensitivity to liquid and ice but rather a bias correction term relying on these variables.
Combining these insights with the Pareto-optimality of the simplified “Wang” scheme (gray star in Fig. 2; Wang et al., 2023, based on the scheme of Xu and Randall, 1996), which outputs:
with in appendix A, suggests that the dependence of cloud cover on condensates is fundamentally non-linear. This justifies the use of deep learning to quickly explore which features can be combined to improve the non-linear, functional representation of cloud cover.
4.1.4 Deep Learning
We start with the baseline model from Grundner et al. (2022), which uses a 3-layer-deep, 64-neuron-wide multi-layer perceptron with batch normalization after the second hidden layer. Hyperparameters were optimized using the SHERPA Python library (Hertel et al., 2020). This NN estimates the target cloud cover with high fidelity (MSE=), but at the cost of increased complexity, as it has a total of 9345 trainable parameters. Note that the models from Grundner et al. (2022) were not designed to minimize the number of trainable parameters, so we do not overly focus on this complexity metric. The MSE can be further lowered to with vertically non-local NNs, which map the entire atmospheric column of inputs to the entire column of outputs without inductive bias. However, this small error gain is deemed insufficient given the increased complexity cost.
Instead, we make the “vertically quasi-local” assumption (Eq. 6) and deploy a hierarchy of Pareto-optimal NNs, with features selected sequentially using cross-validated MSE. The five most informative features for NNs are:
Unlike features selected by polynomial models, these can be non-linearly combined to yield high-quality predictions, as shown in the right panels of Fig. 2. First, NNs improve the functional representation of cloud cover by accurately modeling the sharp increase in cloud cover above a temperature-dependent relative humidity threshold – a highly nonlinear, bivariate behavior that simple schemes struggle to capture. Formally:
where is a Pareto-optimal NN at the bottom-right of the (complexity, error) plane. Second, by incorporating vertical relative humidity gradients, NNs can capture stratocumulus decks off the coasts of regions like California, Peru/Chile, and Namibia/Angola. This improvement is especially visible when comparing the error map of an NN using the first four features from Equation 16 to that of an NN additionally incorporating (Fig 2, bottom-right panel).
While indicative of how accurately cloud cover can be parameterized, such improvements are often insufficient to be considered “discoveries” as they remain hard to explain, even with post-hoc explanation tools (Fig. 8 & 9 of Grundner et al., 2022). Therefore, improvements in functional representation and feature assimilation need to be further distilled into sparse models that scientists can readily interpret.
4.1.5 Symbolic Distillation and Equation Discovery
For this purpose, we use symbolic regression libraries, which optimize both the parameters and structure of an analytical model within a space of expressions. Symbolic regression yields expressions with transparent out-of-distribution behavior (asymptotics, periodicity, etc.) (La Cava et al., 2021), making them well-suited for high-stakes societal applications (Rudin, 2019) and the empirical distillation of natural laws (Schmidt and Lipson, 2009). To avoid overly restricting the analytical form of the distilled equation, genetic programming is used. Genetic programming evolves a population of mathematical expressions using methods such as selection, crossover, and mutation to improve a fitness function (Koza, 1994). Given that genetic programming scales poorly with the number of features (Petersen et al., 2019), our NN feature selection results are used to restrict our features to those listed in Equation 16. Using NN results is appropriate since no assumption is made about the type of equation to be discovered.
The GP-GOMEA (light green crosses in Fig. 2; Virgolin et al., 2021) and PySR (dark green crosses in Fig. 2; Cranmer, 2023) libraries were chosen for their ease of use and high relative performance compared to 12 other recent libraries (La Cava et al., 2021). They yielded over 500 closed-form equations for cloud cover, from which the 9 most physically consistent and lowest-error fits were retained with their outputs always clipped to (see Grundner et al., 2024 for details). By physically interpreting the learned parameters, the output of the Pareto-optimal PySR model may be written as:
where the first term may be interpreted as a sparse, third-order Taylor expansion around the training-mean relative humidity () and temperature ():
Dominant Taylor series expansion terms are expected when discovering closed-form, subgrid-scale parameterizations (Jakhar et al., 2024), but as PySR is based on genetic programming, we find two more surprising terms:
where can be interpreted as a characteristic height for low-cloud humidity gradients, and () is the value of the relative humidity gradient that maximizes cloud cover at the inversion level. The last term is:
which is a monotonically increasing function of the condensates’ concentrations, whose trainable parameters are provided in appendix A. Consistently, the Pareto-optimal GP-GOMEA equation has a term that sharply increases as liquid or ice concentrations exceed 0:
where the trainable parameters are listed in appendix A.
In addition to being easily transferable thanks to their low number of trainable parameters, the added value of these equations is transparent: the improved functional representation is explicit (see Eq. 19), and the assimilation of new features is interpretable (see Eq. 20). Finally, scientific discovery may arise through the unexpected aspects of these equations that are robust across models, such as the difference between how cloud cover reacts to an increase in environmental liquid versus ice content. Indeed, at high resolution, cloud cover will become 1 as soon as condensates exceed a small threshold (here kg/kg) independently of the water’s phase. Then, how come cloud cover is more sensitive to ice than liquid in Eq. 21 () and increases much faster with ice than liquid (for ) in Eq. 23?
These are in fact emerging properties of the subgrid distributions of liquid and ice (Grundner et al., 2024): As large values of cloud ice are rarely observed, larger spatial averages of cloud ice at coarse resolution means that many more high-resolution pixels contain low values of cloud ice compared to the liquid case, resulting in higher cloudiness for a given spatially-averaged condensate value. By assuming an exponential distribution for the subgrid liquid and ice content, we can even interpret and as the rate of the respective exponential distributions. This allows us to hypothesize that the nonlinear relationship between condensates and cloud cover is not scale-invariant and requires separate treatments of liquid and ice, with implications for the interaction between microphysical processes and the radiative budget. While analyzing the feature importance of liquid and ice in neural networks could have suggested this difference, it would have been difficult to fully explain it and bridge spatial scales without distillation, confirming the importance of only treating deep learning as a first and not final step towards knowledge discovery. We now turn to a higher-dimensional problem in which spatial connectivity has to be considered: radiative transfer.
4.2 Shortwave Radiative Transfer Emulation to Accelerate Numerical Weather Prediction
4.2.1 Motivation
The energy transfer resulting from the interaction between electromagnetic radiation and the atmosphere, known as radiative transfer, is costly to simulate accurately. Line-by-line calculations of gaseous absorption at each electromagnetic wavelength (Clough et al., 1992) are too expensive for routine weather and climate models. Instead, models often use the correlated- method (Mlawer et al., 1997), which groups absorption coefficients in a cumulative probability space to speed up radiative transfer calculations without significantly compromising accuracy. However, even the correlated- method imposes a high computational burden (Veerman et al., 2021), forcing most simulations to reduce the temporal and spatial resolution of radiative transfer calculations, which can degrade prediction quality (Morcrette, 2000; Hogan and Bozzo, 2018).
This challenge has driven the development of ML emulators for radiative transfer in numerical weather prediction since the 1990s (Cheruy et al., 1996; Chevallier et al., 1998, 2000). ML architectures have become more sophisticated (e.g., Belochitski and Krasnopolsky, 2021; Kim and Song, 2022; Ukkonen, 2022), but the primary goal remains to emulate the original radiation scheme as faithfully as possible. This allows the reduced inference cost of the ML model, once trained, to be leveraged for running the atmospheric model coupled with the emulator, enabling less expensive and more frequent radiative transfer calculations.
In this section, we examine how ML architectural designs impact the reliability of shortwave radiative transfer (covering solar radiation and wavelengths of 0.23-5.85 m). As shown in Fig. 3, architectures that closely mimic the vertical bidirectionality of radiative transfer are Pareto-optimal, rivaling the performance of deep learning models with ten times more trainable parameters.
4.2.2 Setup
We follow the setup described in Lagerquist et al. (2022) and emulate the full behavior (gas optics and radiative transfer solver) of the shortwave Rapid Radiative Transfer Model (Mlawer et al., 1997) in the context of weather predictions made by version 16 of the Global Forecast System with horizontal spacing. In addition to the realistic geography setup of Lagerquist et al. (2021), the ML models are trained with global data on the 127-level native pressure-sigma grid (henceforth referred to as ) with synthetic information about aerosols, trace gases, and hydrometeors’ particle size distribution. We use data from most days between Sep 1 2018 and Dec 23 2020, with forecast lead times in hours. In each forecast, 4000 grid points are randomly sampled on the global grid. The training set comprises 873,086 samples from 237 days, the validation set 479,806 samples from 126 days, and the test set 472,456 samples from 120 days. Features can broadly be separated into 22 vertical profiles tied to the grid (; listed in Appendix B) and four scalars with one value per atmospheric column: solar zenith angle [∘], surface albedo , aerosol single-scattering albedo , and aerosol asymmetry parameter . For convenience, the four scalars are converted into profiles by repeating their values at all 127 levels. We target the shortwave heating rate’s vertical profile [K day-1], thus approximating a high-dimensional mapping:
via an ML model hierarchy described in the next section.
4.2.3 Machine Learning Model Hierarchy
We deploy an ML model hierarchy using the same features to isolate the added value of spatial connectivity along the vertical coordinate . For all models, cropping and zero-padding layers set the top-of-atmosphere (TOA; 78 km above ground level) heating rate to zero, consistent with the RRTM model being emulated. Readers interested in the exact ML model architecture are referred to Lagerquist et al. (2022) and the accompanying code (https://zenodo.org/records/12557544).
We begin with a linear regression model (pink star), where the input layer is reshaped into a flattened vector and processed through linear layers without activation functions or pooling. This model lacks inherent spatial connectivity, treating all features (where one “feature” = one atmospheric variable at one vertical level) independently. Similarly, the multilayer perceptrons (MLP; orange crosses) flatten inputs, ignoring vertical connectivity, but process them with varying degrees of nonlinearity, controlled by two hyperparameters: depth (number of dense layers) and width (number of neurons per layer). We conduct a grid search with depth varying from layers and width varying from neurons, resulting in 36 MLPs. The activation function for every hidden layer (i.e., every layer except the output) is the leaky rectified linear unit (ReLU; Maas et al., 2013) with a negative slope .
To incorporate vertical relationships between different levels, we train convolutional neural networks (CNN; red crosses) with 1D convolutions along the vertical axis. We conduct a grid search with depth varying from convolutional layers and number of channels in the first convolutional layer varying from . After the first layer, we double the number of channels in each successive convolutional layer, up to a maximum of 64. The CNN’s kernel size is restricted to 3 vertical levels, to enforce strong local connectivity.
The U-net architecture (pink crosses; Ronneberger et al., 2015) builds on the CNN by incorporating skip connections that preserve high-resolution spatial information and an expansive path almost symmetric to the contracting path. Both of these model components improve the reconstruction of full-resolution data (here, a 127-level profile of heating rates). We conduct a grid search with depth (number of downsampling operations) varying from and number of channels in the first convolutional layer varying from , using the same channel-doubling rule as for CNNs. Finally, the U-net++ architecture (purple crosses; Zhou et al., 2019) enhances the U-net’s skip pathways with nested dense convolutional blocks, potentially capturing more of the original spatial information and facilitating optimization. Our hyperparameters for U-net++ are the same as for the U-net.
4.2.4 Distilling Radiative Transfer’s Bidirectionality
To better understand the added value of each architecture in representing shortwave absorption and scattering, we extract samples from our test set to form two distinct regimes. The simple, “clear-sky” regime (Fig. 3a) consists of profiles with no cloud (column-integrated liquid-water path = column-integrated ice-water path = 0 g m-2), an oblique Sun angle (zenith angle 60∘), and little water vapor (column-maximum specific humidity 2 g kg-1). This restricts shortwave radiative transfer to gaseous absorption of solar radiation throughout the atmospheric column, well described by a simple exponential attenuation model such as Beer’s law (Liou, 2002). In contrast, the complex, “multi-cloud” regime (Fig. 3b) includes profiles with more than one cloud layer and at least one mixed-phase cloud layer. For this purpose, a “cloud layer” is defined as a set of consecutive vertical levels, such that every level has a total water content (liquid + ice) 0 g m-3 and the layer has a height-integrated total water path (liquid + ice) 10 g m-2. A mixed-phase cloud layer meets the above criteria plus two additional ones: both the height-integrated liquid water path and ice water path must be 0 g m-2. This regime is challenging to model as shortwave radiation is absorbed and scattered by both liquid and ice clouds, making the bidirectionality of radiative transfer essential to capture. Out of the 472,456 test set profiles, 14,226 are in the clear-sky regime and 13,263 in the multi-cloud regime.
The first surprising result is the poor performance of CNNs: While Pareto-optimal for low complexity, with our simplest CNN having as few as 12,630 trainable parameters, none of the CNNs achieve an MSE below 0.95 K2 day-2, even in the clear-sky case. In contrast, MLPs systematically outperform our linear baseline model, showcasing the importance of nonlinearity. MLPs also outperform CNNs because they allow for non-local vertical connectivity. However, without information on which vertical levels are closest, MLPs connect every level together, resulting in high complexity: Increasing the number of trainable parameters by 100 does not even halve the MSE in the clear-sky case. This lack of inductive bias prevents generalization to more complex cases such as the multi-cloud regime, where the MSE climbs from 0.2-0.4 K2 day-2 to 0.7-1.0 K2 day-2.
By ordering the features’ vertical levels, both the U-net and U-net++ achieve lower errors with fewer trainable parameters. Most Pareto-optimal models are in the U-net++ family, suggesting that simple skip connections are insufficient; this highlights the non-locality of shortwave radiative transfer, especially in the multi-cloud case. Fortunately, while shortwave radiation may instantaneously propagate information (e.g., the presence of mixed-phase clouds) throughout the entire atmospheric column, its transfer is approximately governed by bidirectional equations invariant in space. The two-stream equation holds exactly for the target Rapid Radiative Transfer Model data:
where and are the upward and downward radiative fluxes (W m-2) and the function governs how these fluxes change with the vertical coordinate , allowing us to update fluxes using neighboring fluxes only. As noted in Ukkonen and Chantry (2024), it is the vertical invariance of Eq. 25 that suggests we may treat radiative transfer with a network processing the vertical sequences of and in both directions with the same update rule at every level. Akin to standard atmospheric models, the target heating rate is then calculated by multiplying the vertical gradient of the net flux (downward minus upward) by the ratio of the gravity constant at the Earth’s surface to the specific heat capacity of dry air at constant pressure :
Following Ukkonen (2022), we divide the shortwave radiative fluxes by the incoming TOA downward flux (), to implicitly inform the NN that the Sun is the original energy source. At each vertical level , we then target the resulting ratios ( and ) with a bidirectional RNN. The RNN has two 64-unit gated recurrent unit layers (one forward, one backward), which are concatenated before a dense layer for predicting the fluxes, resulting in only 54,786 trainable parameters. The RNN includes a multiplication layer with and a custom layer to directly calculate shortwave heating rates, maintaining the same optimization objective as the other models. We find that the resulting RNN (green star) is clearly Pareto-optimal, competing with the U-net++ and achieving MSE below 0.1 K2 day-2, even in the complex, multi-cloud case. Our results highlight the advantages of physically constraining ML solutions using robust knowledge of the underlying system, extending the aquaplanet findings of Bertoli et al. (2023) to an Earth-like, operational setting, and paving the way towards fully hybrid physics-ML emulators of shortwave radiation (Schneiderman, 2024). We now turn to a problem in which both spatial and temporal connectivities may be considered: precipitation parameterization.
4.3 Tropical Precipitation and Convective Organization
4.3.1 Motivation
Accurately representing precipitation processes in tropical regions enhances global Earth system models and forecasting tools, particularly for water management and flood risk in a changing climate (Seneviratne et al., 2021). Due to computational limitations, achieving horizontal resolutions below 25-50 km is challenging, which hinders the representation of subgrid processes causing precipitation (Stevens et al., 2020). These processes include tropical convection and its complex organization patterns (Bao et al., 2024), typically modeled using semi-empirical parameterizations that rely on a coarse representation of the thermodynamic environment and often exclude memory effects (Colin et al., 2019). These parameterizations generally approximate a mapping: , where at the time of interest , both the spatially resolved environmental predictors and the precipitation are averaged over the coarse grid cell:
where represents a continuous, bidimensional horizontal coordinate. This spatial averaging removes convective organization information below the coarse grid’s spatial scale. Using storm-resolving simulations as a reference, Shamekh et al. (2023) showed this blurs precipitation, adding irreducible uncertainty, especially for large values of precipitable water (PW). Given that high-resolution information is typically inaccessible in a parameterization context, we ask in this section: How much of this lost spatial granularity can we recover with temporal memory?
4.3.2 Setup
To address this question, we build upon the framework of Shamekh et al. (2023), coarse-graining SAM-DYAMOND (Khairoutdinov et al., 2022) simulations from their native, high-resolution horizontal grid (4.34-km spacing at the equator) to a low-resolution grid (138.9-km spacing at the equator) representative of a coarse Earth system model. Over the tropical ocean (S-N), we map 6 environmental variables – PW, 2-m specific humidity [kg kg-1], 2-m temperature [K], surface sensible and latent heat fluxes [], and sea surface temperature [K] – to 15-min precipitation rates [mm h-1]. We use three setups at a given low-resolution location including the high-resolution grid locations :
where PW′ is PW’s anomaly with respect to its grid cell-mean and memory is accounted for through up to two time steps in the past, in which case . Samples are extracted from 10 days of the simulation after spin-up, with 6 days for training, 2 for validation, and 2 for testing. To focus on precipitation intensity rather than triggering, samples with precipitation values below 0.05 mm/h are discarded, resulting in a total of samples.
4.3.3 Comparing Spatial and Temporal Connectivities
As illustrated in Fig. 4, we design three categories of NNs to learn the three mappings given by Eq. 28, 29, and 30. First, our baseline NNs (orange crosses) are simple MLPs with 5 layers of different widths, resulting in a number of trainable parameters ranging from and never achieving MSEs below 0.4 mm2 h-2. The NNs using spatial granularity (purple crosses) additionally encode the high-resolution PW anomaly field through an encoder-decoder structure with a small bottleneck representing convective organization as a bidimensional latent variable, optionally regularized via data augmentation applied to PW′. MSEs are below 0.04 mm2 h-2 for the deepest encoder-decoder architectures, which yield high reconstruction quality through more than 1M trainable parameters. The NNs leveraging temporal memory instead of spatial resolution (green crosses) use the last or last two previous timesteps to achieve competitive MSEs, all below 0.15 mm2 h-2. These “Memory-NNs” are built using two types of architectures: On the one hand, we flatten previous timesteps and feed them to MLPs conditioned on non-zero current precipitation, which ignores temporal connectivity. On the other hand, we use CNN-based temporal models that preserve temporal ordering through 1D convolutional layers, achieving MSEs around 0.1 mm2 h-2 with times fewer trainable parameters than the “High-Res Inputs-NNs” that use storm-scale information. Overall, our results showcase the added value of spatial and temporal connectivities, as while:
4.3.4 Mitigating Low Spatial Resolution with Memory
While it may be unsurprising that leveraging spatio-temporal information decreases error, the competitive MSEs obtained with temporal memory but without high-resolution spatial information are promising for improving parameterizations. We ask: How can we explain that temporal memory helps recover a large portion of the spatial granularity information?
First, we can ask: Why would temporal memory of the coarse environment be relevant on its own, given that at a given time , the precipitation can be exactly diagnosed from the high-resolution environment ? The informativeness of high spatial resolution is confirmed by the fact that our lowest-error model is . As written in Eq. 29, decomposes the high-resolution information into a low-resolution component and the inputs’ spatial anomaly . Using this decomposition, we can write the following causal graph:
where we posit that for a small enough time step , the combination of the high-resolution anomaly and low-resolution information is enough to diagnose precipitation and step the entire system forward in time from to . Formally, this is a Markovian assumption and eliminates the need for temporal memory. Using Eq. 32 at times and , we see that in this framework, suppressing access to high-resolution information “unblocks” the following causal path:
creating a conditional dependence of precipitation on the past low-resolution environment by removing their d-separation given current high-resolution anomalies (Pearl et al., 2000). This explains the need for temporal memory without high-resolution information and provides a rationale for the Pareto optimality of the “simple-memory NN” that considers the low-resolution inputs and their past timesteps.
Now, we can ask: How much information is lost in the process of replacing the high-resolution spatial anomaly with temporal memory of the low-resolution spatial means? The answer depends on the distance between low and high resolutions. There is no difference if these resolutions are equal, while we do not expect low-resolution memory to help at all if the low-resolution grid is coarse to the point that we cannot distinguish very different high-resolution situations. For the resolutions considered here, the bidimensionality of the encoder-decoder’s bottleneck suggests that the high-resolution anomaly field can be represented by a bidimensional variable for the purpose of modeling precipitation: , which represents the aspects of convective organization that explain why precipitation can be different (stochastic) for the same low-resolution inputs (Moseley et al., 2016). This simplifies the question to how well we can represent from the time series of low-resolution inputs. The encouraging results of our models leveraging temporal memory, along with the finding of Shamekh et al. (2023) that is accurately predicted by an autoregressive model informed by the coarse-scale environment’s past timesteps, suggest a mostly positive response. This reassuringly confirms that Earth system models with limited spatial resolutions can realistically represent coarse-scale precipitation as long as efforts to improve precipitation parameterization continue (Schneider et al., 2024). More broadly, this confirms that Pareto-optimal model hierarchies are useful in empirically establishing the partial equivalence between temporal memory and spatial granularity. This has practical applications for multi-scale dynamical systems that are not self-similar, where ergodicity cannot be used to deterministically infer detailed spatial information from coarse time series data.
5 Conclusion
In this study, we demonstrated that Pareto-optimal model hierarchies within a well defined (complexity, error) plane not only guide the development of data-driven models—from simple equations to sophisticated neural networks—but also promote process understanding. To distill knowledge from these hierarchies, we propose a multi-step approach: (1) use models along the Pareto front to hierarchically investigate the added value (functional representation, feature assimilation, spatial and temporal connectivities) that leads to incremental error decreases from the simplest to the most complex Pareto-optimal model; (2) generate hypotheses and propose models tailored to the system of interest based on this added value; and (3) once the models are sparse enough to be interpretable, reveal the system’s underlying properties that previous theories or models may have overlooked. Beyond knowledge discovery, such hierarchies promote interpretability by explaining the added value of models that may initially seem overly complex, and sustainability by optimizing models’ computational costs once their added value is justified.
We have chosen three weather and climate applications in realistic geography settings to showcase the potential of machine learning in bridging fundamental scientific discovery and increasing operational requirements. Each case demonstrates a different nature of discovery: data-driven equation discovery for cloud cover parameterization, physics-guided architectures, informed by the available Pareto-optimal solution, that reflect the bidirectionality and vertical invariance of shortwave radiative transfer, and spatial information hidden within time series of the coarse environment when diagnosing tropical precipitation. However, all three insights required retaining the full family of trained models and might have been overlooked had we focused on a single optimal model.
In all three cases, neural networks proved particularly advantageous as they can quickly explore large datasets and generate hypotheses about problems that are or appear high-dimensional. We nonetheless emphasize that within this framework, deep learning is an integral part but not the ultimate goal of the knowledge generation process. Simpler models rivaling the accuracy of deep neural networks, initially intractable, may emerge once the necessary functional representations, features, and spatio-temporal connectivities are distilled. From this perspective, a rekindled interest in multi-objective optimization and hierarchical thinking would open the door to extracting new, human-interpretable scientific knowledge from ever-expanding geoscientific data, while paving the way for the adoption of machine learning in operational applications by fostering informed trust in their predictions.
Acknowledgements.
[…]
For the GP-GOMEA equation: , , , , , , , , and (0 cloud cover assigned in the absence of condensates and model calibrated only with condensates present).
[…]
The 22 features with a vertical profile are temperature [K], pressure [Pa], specific humidity [kg kg-1], relative humidity [%], liquid water content [kg m-3], ice water content [kg m-3], downward and upward liquid water path [kg m-2], downward and upward ice water path [kg m-2], downward and upward water vapor path [kg m-2], O3 mixing ratio [kg kg-1], height above ground level [m], height layer thickness [m], pressure layer thickness [Pa], liquid effective radius [m], ice effective radius [m], N2O concentration [ppmv], CH4 concentration [ppmv], CO2 concentration [ppmv], and aerosol extinction [m-1].